OpenStudy (anonymous):

sqrt(-cos(t)+cos(t)+1) = sqrt(2)???

OpenStudy (jamesj):

-cos(t)+cos(t)+1 = 1 hence sqrt(-cos(t)+cos(t)+1) = sqrt(1) = 1

OpenStudy (anonymous):

er my bad 1 sec

OpenStudy (anonymous):

\[\sqrt{-\sin(t) + \cos(t) + 1}\]

OpenStudy (anonymous):

typed wrong thing in

OpenStudy (jamesj):

It's not hard to convince yourself that this is not a constant. For example, when x = 0, this is sqrt(2). But when x = pi/4, it is 1.

OpenStudy (anonymous):

yeah, no that's what i mean. i'm doing arc length and curvature in cal 3, but the book has an example problem that went from that equation = sqrt(2) with no bounds

OpenStudy (jamesj):

When you've figured out your question, post it again.

OpenStudy (anonymous):

here i'll post pic real fast

OpenStudy (anonymous):

and how did it go from 2pi to just pi

OpenStudy (jamesj):

Oh, it's written incorrectly. It should by sqrt( (-cos t)^2 + (sin t)^2 +1 ) = sqrt(1 + 1) = sqrt(2)

OpenStudy (anonymous):

see that's what i was thinking bc it would be the regular distance formula but the book has the formula for this written as.. 1 sec posting

OpenStudy (jamesj):

and it should be the integral 0 to 2pi, as the final answer suggests. But that's a lot of typos in a solution set, and a criminally large number of types for just one problem.

OpenStudy (jamesj):

...number of typos ...

OpenStudy (anonymous):

yeah hate how this book just skips steps without justifying anything

OpenStudy (jamesj):

Also wrong, because it's missing the ^2, the squares of all those derivatives.

OpenStudy (anonymous):

wtf ahah

OpenStudy (anonymous):

that's what i was thinking.. 1 sec gonna double check online real fast

OpenStudy (anonymous):

yeh even patrickjmt on youtube has the squares in it

OpenStudy (jamesj):

I have no idea who/what patrickjmt is, but ok.